# Aligning Human Motion Generation with Human Perceptions

### v1

@@ Appendix C Details on MotionCritic: as Motion Quality Metric
Each multiple-choice question is divided into three ordered preference pairs. Motion sequences are parameterized using SMPL [27], which includes 24 axis-angle rotations and one global root translation.
@@
We train the critic model from scratch using the DSTformer [52] backbone with 3 layers and 8 attention heads on MotionPercept. To ensure robustness, we train our model for multiple times and report the error bars, considering variations such as the random seed across multiple runs. Evaluation results, detailing action-label splits, are presented in the following two tables. Our MotionCritic gets the best results and can robustly score different types of human motions.
@@
Acceleration [45] 66.7 78.0 61.6 53.0 65.4 62.4 82.6 61.6 51.1 59.3 69.5 61.7 64.3
Person-Ground Contact [32] 69.8 70.1 70.2 66.0 72.8 71.8 90.1 76.9 70.3 67.9 62.6 63.2 71.8
Foot-Floor Penetration [32] 47.1 52.4 52.7 55.7 48.5 56.8 59.9 52.4 50.4 55.7 52.8 53.3 53.6
Physical Foot Contact [40] 80.5 77.8 73.1 51.7 67.5 57.9 78.5 63.4 53.2 65.5 68.6 68.5 64.8
MoBERT [41] 67.4 68.4 44.8 37.4 70.0 18.9 43.6 49.9 65.2 25.6 56.1 44.8 49.4
@@
Acceleration [45] 0.70 0.59 0.88 1.5 0.60 0.69 0.60 0.64 0.88 0.71 0.61 0.76 0.78
Person-Ground Contact [32] 0.71 0.68 0.68 0.71 0.69 0.70 0.73 0.68 0.74 0.70 0.73 0.72 0.73
Foot-Floor Penetration [32] 0.70 0.70 0.69 0.70 0.69 0.70 0.71 0.70 0.69 0.70 0.69 0.69 0.69
Physical Foot Contact [40] 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69
MoBERT [41] 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.70 0.69 0.69 0.69
@@
Appendix D Details on MotionCritic: as Training Supervision
D.1 Fine-tuning
@@
Inspired by [44], we observe how the critic score changes over denoising steps to identify the optimal time window for ReFL intercept. As shown in Figure 10(A), we set the hyperparameter step sampling range to $[T_{1},T_{2}]=[700,900]$, where the critic score witnesses a rapid increase.
Figure 10(B) illustrates the variation in the average critic score of a training batch over the course of fine-tuning steps. The fine-tuning process is stable and quick to take effect.
@@
Figure 10: Fine-tuning process. (A): Critic score in 1000-step denoising process. (B): Critic output in 800-step fine-tuning process.
D.2 Results
@@
We conduct an independent user study to compare motion pairs generated at various fine-tuning stages and calculate the Elo Rating [7, 40]. Figure 12 demonstrates that the quality of motions consistently enhance as fine-tuning advances, as indicated by the user study. This improvement aligns with the training objective of elevating the critic score.
We further inspect the change of different metrics during the fine-tuning process in Figure 12(B). PFC [40] and FID are expected to be negatively correlated with motion quality (the smaller, the better), and MotionCritic and multimodality are expected to be positively correlated (the greater, the better). The results indicate that existing motion quality metrics (e.g. FID, PFC) do not adequately reflect human preference, as they poorly correlate with Elo ratings from user studies. Meanwhile, improving the critic score does not necessarily conflict with the multimodality metric, which models the diversity of generated motions.
@@
Figure 12: Results from fine-tuning process. (A): Elo ratings from user study, FID and average Critic scores in the fine-tuning process. (B): FID, PFC[40], Multimodality and Critic scores in the fine-tuning process.
Appendix E Details on User Studies
Annotation
We conduct user studies on GT subsets grouped from HumanAct12 [13] and motions generated during finetune steps as discussed in the main text. Our user study platform is shown in Fig 13. In user study, one motion pair of two motions are played simultaneously, with their critic scores and text prompts being hidden. Annotators should choose the better motion or choose "Almost the Same" if they can’t make a decision. We perform user study on 5 different finetune steps and 5 GT batches grouped from HumanAct12 [13].
Win-rates

### v2

@@ B.1 Implementation Details
Each multiple-choice question is divided into three ordered preference pairs. Motion sequences are parameterized using SMPL (Loper et al., 2015), which includes 24 axis-angle rotations and one global root translation.
@@
We train the critic model from scratch using the DSTformer (Zhu et al., 2023) backbone with 3 layers and 8 attention heads on MotionPercept. To ensure robustness, we train our model for multiple times and report the error bars, considering variations such as the random seed across multiple runs. Evaluation results, detailing action-label splits, are presented in the following two tables. Our MotionCritic gets the best results and can robustly score different types of human motions.
@@
Acceleration (Yang et al., 2023) 66.7 78.0 61.6 53.0 65.4 62.4 82.6 61.6 51.1 59.3 69.5 61.7 64.3
Person-Ground Contact (Rempe et al., 2021) 69.8 70.1 70.2 66.0 72.8 71.8 90.1 76.9 70.3 67.9 62.6 63.2 71.8
Foot-Floor Penetration (Rempe et al., 2021) 47.1 52.4 52.7 55.7 48.5 56.8 59.9 52.4 50.4 55.7 52.8 53.3 53.6
Physical Foot Contact (Tseng et al., 2023) 80.5 77.8 73.1 51.7 67.5 57.9 78.5 63.4 53.2 65.5 68.6 68.5 64.8
MoBERT (Voas et al., 2023) 67.4 68.4 44.8 37.4 70.0 18.9 43.6 49.9 65.2 25.6 56.1 44.8 49.4
@@
Acceleration (Yang et al., 2023) 0.70 0.59 0.88 1.5 0.60 0.69 0.60 0.64 0.88 0.71 0.61 0.76 0.78
Person-Ground Contact (Rempe et al., 2021) 0.71 0.68 0.68 0.71 0.69 0.70 0.73 0.68 0.74 0.70 0.73 0.72 0.73
Foot-Floor Penetration (Rempe et al., 2021) 0.70 0.70 0.69 0.70 0.69 0.70 0.71 0.70 0.69 0.70 0.69 0.69 0.69
Physical Foot Contact (Tseng et al., 2023) 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69
MoBERT (Voas et al., 2023) 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.69 0.70 0.69 0.69 0.69
@@
Dataset Splitting and Model Retraining.
MotionPercept training and testing split is randomly divided in the paper. Here we also investigate into division based on annotator IDs. By adopting this approach, we ensure that each annotator’s data is used exclusively in one of the sets, preventing overlap. Subsequently, we retrained and evaluated the model using this revised split. Disalignment between annotators as shown in Figure 2 results in slight decline of MotionCritic performance. Results are shown in the following table and align with our initial expectations.
Metric MDM FLAME
Acc. (%) $\uparrow$ Log Loss $\downarrow$ Acc. (%) $\uparrow$ Log Loss $\downarrow$
Root AVE 59.22 0.6875 48.42 0.6984
Root AE 62.10 0.6787 59.54 0.6711
Joint AVE 57.01 0.6875 44.61 0.6973
Joint AE 63.17 0.6790 58.37 0.6891
Jerk 66.33 0.7476 65.84 0.5984
Acceleration (Yang et al., 2023) 63.82 0.7641 66.67 0.6919
Person-Ground Contact (Rempe et al., 2021) 68.08 0.6835 69.82 0.7243
Foot-Floor Penetration (Rempe et al., 2021) 51.77 0.6972 55.56 0.6906
Physical Foot Contact (Tseng et al., 2023) 66.52 0.6927 66.00 0.6930
PoseNDF (Tiwari et al., 2022) 56.00 0.6930 53.07 0.6931
NPSS (Gopalakrishnan et al., 2019) 53.48 0.6975 52.07 0.6944
NDMS (Tanke et al., 2021) 62.08 0.6590 63.35 0.6301
MoBERT (Voas et al., 2023) 47.49 0.6933 52.40 0.6932
MotionCritic (Ours) 80.13 0.5402 79.24 0.5927
Table 7: Quantitative comparison of metrics on re-split testsets of MotionPercept.
Inplementation Details of Other Metrics
Here we provide details of the implementation of the compared metrics. To ensure reproducibility, we will publish the code for all metrics. Overall, we utilized official code and models whenever possible, making only minimal and reasonable modifications to meet our evaluation needs.
  • Root AVE, Root AE, Joint AVE, and Joint AE: These metrics were implemented following the definitions provided by Voas et al. (2023).
  • Acceleration and Jerk: We used the released official implementations and hyperparameters from Yang et al. (2023) without any modifications. Jerk was derived as the first derivative of Acceleration.
  • Person-Ground Contact and Foot-Floor Penetration: We used the released official implementations and hyperparameters from Rempe et al. (2021) without any modifications.
  • Physical Foot Contact: We followed the approach described in Tseng et al. (2023), with slight adaptations to rewrite the original NumPy implementation into PyTorch.
  • PoseNDF: For each motion sequence, we extracted all frames as poses and calculated neural distances using the official pretrained models provided by Tiwari et al. (2022). We implemented two methods: one using the average distances across all frames as the motion evaluation metric, and the other using the maximum distance among all frames. The latter performed better and was reported in Table 1.
  • NPSS: We used the released official implementations and hyperparameters from Karras et al. (2019) without any modifications. Motions were matched to corresponding ground-truth sequences in the HumanACT12 action class, as NPSS is uni-modal and compares motions to single ground-truth sequences.
  • NDMS: We used the released official implementations and hyperparameters from Tanke et al. (2021) without any modifications. We first transformed axis-angle SMPL into xyz coordinates and utilized code from official implementation for skeletons.
  • MoBERT: We utilized their pretrained model from Voas et al. (2023) and computed the Naturalness metric, which evaluates the intrinsic plausibility of the motion itself and aligns with the theme of our work.
This setup ensures that our evaluation metrics are consistent with the official implementations and provides a clear methodology for reproducibility.
B.2 Failure Modes
In addition, we analyze the failure modes of MotionCritic when it is used as a motion quality metric (Table 8). To do this, we computed the critic score differences for all samples in the test set (user-annotated better vs. user-annotated worse) and summarized the statistics in the table below:
Metric Mean Std Max Min
All Scores 4.32 3.72 29.82 -11.73
Correct Cases (85.1%) 4.75 3.79 29.82 6.79e-5
Wrong Cases (14.9%) -1.84 1.91 -1.91e-4 -11.73
Table 8: Critic score statistics for all cases, correct cases, and wrong cases.
We note that in correct cases, the critic score differences are larger, indicating higher confidence. Conversely, in the wrong cases, the score differences are smaller, suggesting that the errors occur when the model is less confident. We also provide the distributions of critic scores for both correct and wrong predictions (see Fig. 10). On the left, the vertical axis represents the number of examples (with wrong cases comprising 14.9% and correct cases 85.1%). On the right, the vertical axis shows the percentage within the corresponding part.
Figure 10: Distributions of critic scores’ differences for correct and wrong predictions.
It is clear that in cases where the model predicts incorrectly, the majority of the score differences are very small. For example, 35.10% of the wrong cases have a critic score difference within 0.73 (one bin), whereas only 8.74% of correct cases fall into the same range. This indicates that many of the incorrect predictions occur when the critic scores are close, rather than due to large, obvious errors. In contrast, correct predictions tend to have larger score differences, reflecting higher confidence from the model. Additionally, some errors occur in particularly challenging examples where even human annotators struggle to differentiate, and different individuals may make varying judgments.
B.3 Sensitivity Analysis
We conducted a preliminary sensitivity analysis by perturbing AMASS ground-truth (GT) motions with Gaussian noise of various scales. Assuming that GT motions should consistently receive higher scores than perturbed ones, we evaluated the critic’s ability to distinguish between them.
Robustness.
MotionCritic demonstrates robustness in ranking GT motions above perturbed ones across noise scales, as demonstrated in Figure 11(A). We also observed how the average and standard deviation of critic scores vary with different noise scales, as demonstrated in Figure 11(B).
Smoothness.
By analyzing the 3D relationship between noise scale, critic scores, and perturbed critic scores, we observed steep score changes near the natural motion distribution, as demonstrated in Figure 12(B). The interactive version of this 3D visualization is available at https://iclrauthors6688.github.io/iclr6688/sensitivity_analysis_3d.html , highlighting the critic’s sensitivity outside the natural motion distribution. We also provide a 2D scatter plot focusing on perturbed-score vs noise-scale, as demonstrated in Figure 12(A).
Non-smooth energy landscape.
In summary, we discover that the critic model exhibits a non-smooth and "bumpy" energy landscape, which serves as a double-edged sword in our framework. On one hand, this non-smoothness is a desirable property, making the critic highly discriminative and sensitive to subtle motion perturbations, aligning well with human perceptual abilities. On the other hand, it introduces challenges to motion optimization. Future work could investigate better optimization methods, such as leveraging reinforcement learning strategies (Liu et al., 2024), to mitigate challenges associated with this non-smoothness.
Non-distance measure.
Based on the training objective, the Bradley-Terry model, the critic scores are intended to represent relative preferences rather than absolute distances. Specifically, the scores indicate whether one motion is judged as superior to another, rather than their distance. Therefore, it’s worth noticing that the critic score should not be interpreted as a distance function. This design aligns with the goal of producing a ranking-based evaluation metric rather than a continuous distance measure.
((A))
((B))
Figure 11: Sensitivity analysis results. (A) Accuracy vs noise-scale curve. (B) Average and standard deviation of critic scores vs noise-scale.
((A)) 2D scatter plot of perturbed-score vs noise-scale.
((B)) 3D scatter plot of noise scale, critic scores, and perturbed critic scores.
Figure 12: Sensitivity analysis upon AMASS GT motions. Critic score differences (original - perturbed) are demonstrated by colors.
B.4 Test on out-of-distribution motions
We test our method on out-of-diftribution motion dataset apart from HumanACT12 and UESTC. Specifically, we conduct additional analyses on AMASS using MotionCritic and observed some interesting findings:
Dataset Mean Std Max Min
AMASS -1.10 4.14 20.77 -18.99
HumanAct12 -2.22 4.68 12.14 -21.96
MDM -2.08 4.65 12.83 -30.65
Table 9: Critic score statistics of the two datasets.
  • The critic scores on AMASS are generally higher than those on HumanAct12, which aligns with our expectations, while the critic score statistics of the two datasets are shown in Table 9.
  • The actions with the lowest scores on AMASS were indeed outliers, often violating physical constraints such as gravity (mostly due to object interactions that appear unreasonable when considering only the motion). At the same time, it was also able to assign relatively high scores to some previously unseen complex motions.
Current results show that AMASS has higher average scores, maximum scores, and minimum scores compared to MDM-generated motions, even though some outliers in AMASS significantly lowered the scores, as discussed above. This suggests that normal motions in AMASS generally outperform those generated by MDM. Additionally, we conduct a pairwise comparison between MDM-generated motions and HumanAct12 GT motions, testing both the critic scores and user study results, which are presented in the table below. HumanAct12 indeed has quality issues , and both the critic score and users often favored the MDM-generated motions. Furthermore, the critic score continues to align well with human users’ judgments when comparing these two types of motions, demonstrating its effectiveness.
Table 10: User study on HumanAct12 GT and MDM-generated motions
Motion Category (12 categories, 10 motion pairs each) HumaAct12 Preferred MDM Preferred User Similar Humact12 Critic MDM Critic User Disagreement
0 10% 70% 20% 20% 80% 20%
1 30% 40% 30% 30% 70% 20%
2 30% 50% 20% 20% 80% 20%
3 20% 50% 30% 20% 80% 10%
4 0% 80% 20% 10% 90% 0%
5 0% 90% 10% 0% 100% 0%
6 40% 10% 50% 70% 30% 0%
7 0% 90% 10% 10% 90% 0%
8 10% 90% 0% 40% 60% 30%
9 10% 80% 10% 50% 50% 30%
10 0% 60% 40% 10% 90% 0%
11 40% 40% 20% 30% 70% 20%
Overall (120 pairs) 15.83% 62.50% 21.67% 25.83% 74.17% 12.50%
  1. 1.
    HumanAct12 User Preference: This column shows the percentage of cases where users preferred the HumanAct12 motions over MDM.
  2. 2.
    MDM User Preference: This column shows the percentage of cases where users preferred the MDM motions over the HumanAct12 motions.
  3. 3.
    User Similar: This column indicates the percentage of cases where users found it difficult to distinguish between the HumanAct12 and MDM motions.
  4. 4.
    HumanAct12 Critic Score Win: This shows the percentage of cases where the HumanAct12 motion received a higher critic score.
  5. 5.
    MDM Critic Score Win: This shows the percentage of cases where the MDM motion received a higher critic score.
  6. 6.
    User Disagreement with Critic Score: This column indicates the percentage of cases where the user’s choice did not align with the critic score.
These examples demonstrate that MotionCritic is an effective tool for automatically comparing and evaluating motion quality, whether individually or in bulk. It also helps uncover valuable insights, further highlighting its utility and potential.
B.5 Impact of sequence length, framerate, and interpolation
To further address concerns regarding the impact of sequence length, framerate, and interpolation on the performance of the MotionCritic model, we conduct a detailed analysis using the AMASS GT dataset.
Dataset and Pre-processing.
We extracted a total of 6,346 OOD (Out-of-Distribution) motion examples from the AMASS GT dataset. Among these, 547 examples (approximately 8.62%) have lengths shorter than 60 frames. We analyzed longer and shorter OOD motions separately to evaluate the model’s performance under varying conditions. To further assess the robustness of the MotionCritic model, we generated a pseudo-GT dataset by adding Gaussian noise at different scales to the motions. We formed better-worse pairs by comparing unnoised and noised versions of the same motion.
For longer motions, we experimented with three different pre-processing approaches:
  1. 1.
    Unprocessed Motion: The full-length motion was fed directly into the model to obtain critic scores.
  2. 2.
    Cut-down to 60 Frames: The motion was truncated to 60 frames before being input into the model, expanding the dataset to a total of 12,011 motions.
  3. 3.
    Uniform Frame Extraction: 60 frames were uniformly extracted from the complete motion, which was then fed into the model for critic scores.
The results, presented as shown in Figure 13, demonstrate that the Critic model performs optimally when the motion lengths are 60 frames, consistent with the original dataset.
Figure 13: Performance of the Critic model across different motion lengths and interpolations.
For shorter motions, two pre-processing approaches were evaluated:
  1. 1.
    Unprocessed Motion: The full-length, short motion was input directly into the model for critic scores.
  2. 2.
    Interpolation to 60 Frames: The motion was uniformly interpolated to 60 frames before being input into the model.
The results, also presented as accuracy-noise line plots (on the left, Fig (A)), indicate that the Critic model can effectively handle shorter sequences with minimal performance loss after interpolation.
The results demonstrate that MotionCritic works best when the motion sequences are 60 frames in length, consistent with the training dataset, which could be achieved through interpolation or extraction. The model also exhibits strong performance across variable lengths and framerates, with only minor reductions in accuracy. This robustness highlights the advantages of the MotionCritic model.
B.6 Ensemble Learning using Different Metrics
To further explore the interrelationship between the different motion quality metrics, we treat the different metrics as features and train a simple classifier to perform binary classification (ensemble learning). To train the ensemble model, we create a separate train-test split from the MotionPercept MDM test set. We explore three ensemble approaches to combine the metrics and evaluate the complementary effects between the heuristics and MotionCritic:
Metric Accuracy (%)
MotionCritic 85.07
Acceleration 64.26
Jerk 65.48
GroundContact 71.78
PFC 64.79
Table 11: Initial accuracy of metrics on the MotionPercept MDM test set.
Method Accuracy (%)
Logistic Regression 81.12
SVM 82.49
MLP 85.17
MotionCritic Only 84.96
Table 12: Accuracy of ensemble learning methods.
  1. 1.
    Logistic Regression. Binary classification using the five normalized features.
  2. 2.
    Support Vector Machine (SVM). Binary classification using the five features across two motions in a better-worse pair.
  3. 3.
    Multilayer Perceptron (MLP). A simple MLP with one hidden layer of 128 dimensions was trained using the five features as input, producing a single score to determine better or worse outcomes. The training objective was based on the Bradley-Terry model, similar to MotionCritic.
The accuracy of each combined method is presented in Table 12. The results indicate that combining the metrics does not significantly improve accuracy beyond what MotionCritic alone can achieve, suggesting that MotionCritic already incorporates much of the heuristic knowledge.
Visualization of Heuristics and MotionCritic Overlap
We visualized the overlap between heuristics and MotionCritic on the entire MotionPercept MDM test set using a Venn diagram at Figure 14, demonstrating a significant overlap in correctly classified pairs. This suggests that when MotionCritic makes a mistake, other metrics tend to err in the same cases.
Ablation Study in SVM
To further understand the contribution of each metric, we conducted an ablation study using SVM by incrementally adding features in the order of [MotionCritic, Acceleration, Jerk, GroundContact, PFC]. The results are shown in Table 13.
Features Added Accuracy (%)
MotionCritic 82.49
MotionCritic + Acceleration 82.40
MotionCritic + Acceleration + Jerk 82.45
MotionCritic + Acceleration + Jerk + GroundContact 82.49
All Features 82.49
Table 13: Ablation study results in SVM.
Feature Removal in SVM
We also conducted experiments by sequentially removing features to observe the impact on accuracy, as presented in Table 15. The results indicate that removing MotionCritic has a substantial negative impact on model performance, while removing other features has a minimal effect.
Logistic Regression Coefficients
The coefficients from the logistic regression model are presented in Table 15. The positive coefficient for MotionCritic indicates that higher values are associated with better outcomes, while the negative coefficients for other metrics suggest that lower values are preferable.
Feature Removed Accuracy (%)
MotionCritic 66.52
Acceleration 82.49
Jerk 82.40
GroundContact 82.40
PFC 82.49
Table 14: Accuracy after feature removal in SVM.
Feature Coefficient
MotionCritic 2.49
Acceleration -0.61
Jerk -0.33
GroundContact -0.39
PFC -0.07
Table 15: Logistic regression coefficients.
SHAP Values Visualization
We visualized the SHAP (SHapley Additive exPlanations) values, which provide insights into the contribution of each feature to the model’s predictions (see Figure 14). The figures show the SHAP values for the five metrics tested on a series of samples from our test set, evaluated with the MLP ensemble model. Across these samples, the SHAP values for the MotionCritic feature generally have the largest absolute values, indicating that the critic score is the most important feature. Additionally, when the critic score is high, its SHAP value is positive, suggesting a positive correlation between the critic score and motion quality, which aligns with expectations.
Figure 14: Metrics’ venn fiagram and SHAP values.
The experimental results demonstrate that MotionCritic effectively captures the essential knowledge provided by the heuristics. Attempts to further enhance accuracy by combining MotionCritic with other heuristics yield marginal improvements, underscoring the robustness of MotionCritic in this context.
Appendix C Details on MotionCritic: as Training Supervision
C.1 Fine-tuning
@@
Inspired by ImageReward(Xu et al., 2024), we observe how the critic score changes over denoising steps to identify the optimal time window for ReFL intercept. As shown in Figure 15(A), we set the hyperparameter step sampling range to $[T_{1},T_{2}]=[700,900]$, where the critic score witnesses a rapid increase. The full fine-tuning curve, provided in Figure 15(C), demonstrates that the critic output improves rapidly during the early stages of fine-tuning and stabilizes after a certain point.
In addition, another discussion by Clark et al. (2024) proposed DRaFT-LV, employing single-step gradient backpropagation, observed to be the most efficient. To evaluate the impact of alternative step sampling strategies, we conducted experiments comparing our current setup with DRaFT-LV’s methodology. The results, detailed in Figure 15(C), show that single-step refinement can indeed improve fine-tuning efficiency, consistent with existing findings.
@@
((C))
Figure 15: Fine-tuning process. (A): Critic score in 1000-step denoising process. (B): Critic output in 800-step fine-tuning process. (C): Full finetuning process of our strategy based on ReFL (Xu et al., 2024) and 1-step back-propagation based on DRaFT-LV (Clark et al., 2024).
C.2 Finetuned Results
@@
We conduct an independent user study to compare motion pairs generated at various fine-tuning stages and calculate the Elo Rating (Elo, 1978; Tseng et al., 2023). Figure 17 demonstrates that the quality of motions consistently enhance as fine-tuning advances, as indicated by the user study. This improvement aligns with the training objective of elevating the critic score.
@@
Figure 17: Results from fine-tuning process. (A): Elo ratings and Critic scores. (B): FID, PFC(Tseng et al., 2023), Multimodality and Critic scores.
We further inspect the change of different metrics during the fine-tuning process in Figure 17(B) and Table 2. PFC (Tseng et al., 2023) and FID are expected to be negatively correlated with motion quality (the smaller, the better), and MotionCritic and multimodality are expected to be positively correlated (the greater, the better). The results indicate that existing motion quality metrics (e.g. FID, PFC) do not adequately reflect human preference, as they poorly correlate with Elo ratings from user studies. Meanwhile, improving the critic score does not necessarily conflict with the multimodality metric, which models the diversity of generated motions.
Appendix D Details on User Studies
D.1 Annotation Details
The annotators of MotionPercept are recruited by an annotation company and are ordinary adults (5 men and 5 women). Previous psychological research has shown that the ability to perceive and discern biological motion is universal and consistent, with a neural basis  (Johansson, 1973), and does not rely on specific training or professional background. This finding is also consistent with what we observed in Figure 2. As for potential biases, while some degree of annotator bias may exist (as with any human-annotated dataset), we believe it is minor and manageable, as indicated by our analysis in Section 3. Overall, we consider the opinions of the annotators to be representative.
D.2 Stand-alone User Studies
Details.
We conduct user studies on GT subsets grouped from HumanAct12 (Guo et al., 2020) and motions generated during finetune steps as discussed in the main text. Our user study platform is shown in Fig 18. In user study, one motion pair of two motions are played simultaneously, with their critic scores and text prompts being hidden. Annotators should choose the better motion or choose "Almost the Same" if they can’t make a decision. We perform user study on 5 different finetune steps and 5 GT batches grouped from HumanAct12 (Guo et al., 2020).
Different annotators.
For the stand-alone user study, we did not use large-scale annotators from the annotation company. Instead, we recruit new volunteers through a campus announcement who performed pairwise comparisons of two motions, selecting the better one. The five volunteers are students from two different universities with backgrounds in computer science, engineering, arts, and history. During this process, they do not see any additional information such as scores or text. We ensure that they receive reasonable compensation for their time and effort. We then calculate the Elo Rating based on their results. The annotation interface and rating formulas can be found in this section as well.
Figure 18: The interface of our user study platform.
D.3 Rating Formulas
Win-rates.